A company with a share price of $57 has shares outstanding of 4 million and net income of $11 million. The company conducts a leverage recapitalization to alter its capital structure by borrowing $4 million and using the proceeds to buy back shares. The pre-tax cost of debt is 6% and the company's tax rate is 39%.

What will be the new earnings per share after the leverage recapitalization?

Enter your answer rounded to two decimals.

Please double check your answer please as several answered it incorrectly. Thanks so much!
Answer #1

shares repurchased=(4*1000000)/57=70175 shares

new shares post shares repurchased=(4*1000000)-70175=3929825

net income after recapitalization=(11*1000000)-(4*1000000)*6%*(1-39%)=10853600

What will be the new earnings per share after the leverage recapitalization=10853600/3929825=2.76 (answer)